### v1

# Lightweight Deep Learning-Based Channel Estimation for RIS-Aided Extremely Large-Scale MIMO Systems on Resource-Limited Edge Devices

Abstract
Next-generation wireless technologies such as 6G aim to meet demanding requirements such as ultra-high data rates, low latency, and enhanced connectivity. Extremely Large-Scale MIMO (XL-MIMO) and Reconfigurable Intelligent Surface (RIS) are key enablers, with XL-MIMO boosting spectral and energy efficiency through numerous antennas, and RIS offering dynamic control over the wireless environment via passive reflective elements. However, realizing their full potential depends on accurate Channel State Information (CSI). Recent advances in deep learning have facilitated efficient cascaded channel estimation. However, the scalability and practical deployment of existing estimation models in XL-MIMO systems remain limited. The growing number of antennas and RIS elements introduces a significant barrier to real-time and efficient channel estimation, drastically increasing data volume, escalating computational complexity, requiring advanced hardware, and resulting in substantial energy consumption. To address these challenges, we propose a lightweight deep learning framework for efficient cascaded channel estimation in XL-MIMO systems, designed to minimize computational complexity and make it suitable for deployment on resource-constrained edge devices. Using spatial correlations in the channel, we introduce a patch-based training mechanism that reduces the dimensionality of input to patch-level representations while preserving essential information, allowing scalable training for large-scale systems. Simulation results under diverse conditions demonstrate that our framework significantly improves estimation accuracy and reduces computational complexity, regardless of the increasing number of antennas and RIS elements in XL-MIMO systems.
Index Terms:
Reconfigurable Intelligent Surface, B6G, Deep Learning, UNet 3+, Channel Estimation, Extremely Large-Scale MIMO, Spatial Correlation in Wireless Channels.
I Introduction
Next-generation communication technologies, such as 6G and beyond, are expected to meet the rapidly growing demand for advanced service requirements [1]. These include ultra-high data rates, extremely low latency, minimal bit error rates, enhanced device connectivity, and much more. To address these challenges, Extremely Large-Scale MIMO (XL-MIMO) has emerged as a transformative paradigm, offering significantly higher spectral and energy efficiency while ensuring reliable communication [2]. Unlike conventional MIMO systems, XL-MIMO integrates thousands of antennas in a compact space, fundamentally altering electromagnetic characteristics [3]. This not only expands the number of antennas, but also enhances spectral efficiency and spatial degrees of freedom, enabling unprecedented improvements in wireless communication performance [4]. Building on existing advantages, the integration of Reconfigurable Intelligent Surfaces (RIS) with XL-MIMO presents a promising approach to address the growing demands of 6G and beyond networks [1].
RIS technology is emerging as a transformative enabler, offering the ability to dynamically shape wireless channels between users and the Base Station (BS), thereby enhancing the overall performance of the system [5]. RIS panels consist of tunable electromagnetic elements, each independently controlled to direct incident waves towards targeted directions [6, 7]. This adaptability optimizes the propagation environment, significantly boosting the spectral efficiency [8]. Furthermore, RIS elements are energy efficient and passive, seamlessly integrated into surfaces such as walls and ceilings, making them a low-energy, cost-effective solution for significantly improving signal quality and network coverage [9].
Realizing the performance gains of RIS-enabled XL-MIMO systems in practice requires optimizing the design parameters, which are highly dependent on accurate Channel State Information (CSI) [10]. However, obtaining CSI in RIS is challenging because RIS elements are passive and do not receive or transmit signals directly [11]. Since individual channels between user–RIS and RIS–BS cannot be estimated separately, only the cascaded channel can be inferred, which encompasses the link from the user to the BS via the RIS. This limitation presents distinct challenges in accurately estimating the overall channel. Firstly, unlike traditional Rayleigh fading channels, RIS-based channels deviate from conventional models, resulting in a performance gap between Minimum Mean Square Error (MMSE) and Linear-MMSE (LMMSE) estimators [1]. Moreover, both RIS and antenna arrays consist of numerous elements, which complicates cascaded channel estimation and significantly expands the parameter space, making channel estimation resource-intensive for edge devices [12, 6]. Consequently, these challenges limit system performance and hinder the large-scale deployment of RIS-assisted XL-MIMO systems. This paper proposes a power-efficient data-driven methodology for cascaded channel estimation at the edge devices for RIS-enabled XL-MIMO systems.
Recent work in cascaded channel estimation can be broadly categorized into two main sets of approaches: conventional model-driven methods [6, 9, 10] and data-driven techniques [11, 12, 13, 14, 15, 16]. Conventional model-driven methods frame cascaded channel estimation as an optimization problem, solved iteratively to improve accuracy. Although these methods have achieved notable advancements, the need for multiple iterative steps in real-time scenarios makes them time-consuming, which is impractical for future low-latency communications [11]. In contrast, data-driven approaches (such as machine and deep learning) are gaining popularity. These techniques offer minimal inference time, making them well suited for real-world, time-sensitive scenarios. Although data-driven methods require relevant training data, they can significantly enhance the efficiency and accuracy of cascaded channel estimation once trained [15].
Recent advances in deep learning have substantially contributed to RIS-based communication systems, delivering robust performance across diverse environmental conditions and enabling low-complexity inference. Various deep learning approaches have been proposed to improve cascaded channel estimation for massive MIMO systems, significantly improving overall system performance [15, 11, 12, 13, 14, 16]. However, these techniques face scalability challenges and are typically limited to massive MIMO systems, where only tens of antennas and BS elements are considered. Extending the training setup to RIS-enabled XL-MIMO systems, involving hundreds of RIS elements and thousands of antennas as proposed in these models, would not be a viable solution due to various reasons. Firstly, these techniques use tens of thousands of samples to train their models. Processing and storing such large datasets in RIS-assisted XL-MIMO-based systems can lead to inefficiencies, introduce substantial computational and memory demands, and increase both training time and resource consumption [11]. Furthermore, training these models on large datasets requires substantial energy resources, making them not only energy inefficient but also raising environmental concerns due to increasing energy consumption. Additionally, the need for highly specialized high-performance hardware to handle the training of such expensive systems adds considerable costs, making it financially challenging to scale these solutions. Finally, with edge computing gaining attention in recent years as a key enabler for distributed low-latency applications, the deployment of these complex models in edge environments becomes a significant obstacle [17]. The limited computational resources and energy constraints at the edge create a major barrier, hindering the practical implementation of these techniques in future communication networks.
Addressing the aforementioned challenges requires a compact and efficient representation of the cascaded channel. We therefore exploit the spatial correlation present in the cascaded channel, as discussed in [18, 19, 20], to enable computationally efficient model training suitable for edge devices in resource-limited settings. To take advantage of this property, we introduce a patch-based training framework that leverages the inherent spatial correlations in the cascaded channel. In this framework, small patches are extracted from each data point to encapsulate the essential features required for accurate channel estimation. This patch-wise processing significantly reduces input dimensionality and overall system complexity, thereby enabling scalable and computationally efficient training, particularly suitable for XL-MIMO systems deployed at the edge. Furthermore, employing efficient deep learning techniques, such as autoencoders, can transform these representative patches into compact latent representations, thus facilitating more efficient learning of spatial dependencies [21]. These latent features effectively preserve and highlight spatial dependencies between RIS elements and BS antennas, facilitating the learning of inherent correlations while substantially reducing overall training complexity. Furthermore, leveraging autoencoders also helps break the correlation of noise in the data, effectively isolating meaningful patterns from unwanted distortions [22].
Emerging autoencoder models, such as UNet, have shown remarkable success in noisy data segmentation and reconstruction by capturing underlying data structures, even in data-limited settings [21]. Unlike conventional autoencoders, UNet enables more precise feature extraction and better performance in complex tasks such as channel denoising. For example, the UNet model has demonstrated strong potential for channel estimation in Massive MIMO systems [23]. Extending the capabilities of the UNet, the UNet3+ architecture leverages multi-scale feature aggregation to better extract relevant structures from noisy inputs, improving denoising performance [24]. Therefore, we adopt the UNet 3+ architecture in our proposed channel estimation framework, leveraging its ability to effectively capture spatial features and improve estimation accuracy, especially in resource-constrained environments such as edge devices. Overall, this is the first work to propose an efficient solution for channel estimation in RIS-assisted XL-MIMO systems. Moreover, we formulate the channel estimation task as a denoising problem, where the goal is to recover clean channel information from noisy observations. The proposed method significantly improves the accuracy of the estimation and reduces computational complexity, regardless of the increasing number of antennas and RIS elements in XL-MIMO systems. Extensive simulations under diverse environmental conditions validate the scalability and effectiveness of the proposed framework, demonstrating substantial gains in cascaded channel estimation performance and overall system efficiency for RIS enabled XL-MIMO deployments.
I-A Literature Review
Efficient cascaded channel estimation has attracted significant attention in recent years due to the promising benefits of RIS. Existing research efforts can be broadly categorized into two main approaches: traditional model-driven methods and data-driven techniques. Traditional model-based approaches for cascaded channel estimation remain widely used due to their ability to operate without training data, allowing direct deployment without prior model training. In [6], the authors tackled the cascaded channel estimation problem by exploiting the sparsity and structural correlations in mmWave environments. Their primary optimizations focused on adjusting the phase shifts of RIS elements and minimizing the pilot overhead required for successive channel updates. The simulation results confirmed the effectiveness of their approach, demonstrating a substantial reduction in training overhead. Moreover, in [7], a binary reflection-based channel estimation approach is proposed, wherein a single element is activated at a time while others remain off, and channel estimates are obtained using the least squares method.
Furthermore, in [9], the authors formulated the cascaded channel estimation task using an alternating optimization algorithm combined with a covariance method, which they solved through a successive convex approximation algorithm. Their findings demonstrated enhanced channel estimation accuracy while reducing pilot overhead. The authors in [10] proposed a channel estimation protocol that leveraged compressive sensing techniques to estimate the cascaded channel. Their approach reduced training overhead by formulating the estimation task as a sparse recovery problem. They introduced a two-step joint channel estimation framework to further improve the estimation performance that exploited the block-sparsity structure shared among multiple users. Despite the adequate performance of these approaches  [6, 9, 10], they rely on numerous iterative steps to achieve optimal results, which limits their efficiency in real-time, low-latency applications when applied to XL-MIMO systems involving hundreds of RIS elements and thousands of antennas. As a result, model-based approaches become impractical for such large-scale deployments due to their excessive complexity and resource demands.
The pioneering work that introduced deep learning for cascaded channel estimation was presented in [15]. The proposed framework consisted of two convolutional neural networks (CNNs), each comprising nine layers, designed to estimate both the direct channel between the BS and the user and the cascaded channel between the BS and the user through an RIS. The simulation results validated the effectiveness of the proposed model, demonstrating improvements over traditional methods such as Least Squares (LS) and multilayer perceptron. Moreover, in [11], the authors proposed a deep residual network (CDRN) method using a CNN-based architecture to effectively remove residual noise from noisy pilot signals, ultimately enhancing estimation accuracy. Additionally, [12] introduced a bidirectional long short-term memory (Bi-LSTM) approach for channel estimation. Bi-LSTM processed past and future information simultaneously through two LSTM layers, one forward and one backward, thereby enhancing feature extraction and achieving superior accuracy in channel estimation compared to other recurrent neural networks.
Similarly, the work in [13] interprets the received pilot signal as a low-resolution image and formulates channel estimation as a super-resolution reconstruction problem. This approach employed a CNN to extract detailed features, followed by a denoising CNN to eliminate additive noise. The simulation results validated the effectiveness of this scheme. Another study in [16] proposed a model-driven deep unfolding approach for cascaded channel estimation. The proposed scheme leverages the rank deficiency of the cascaded channel to achieve accurate estimation while significantly reducing communication overhead. In [14], a self-supervised learning approach was proposed for channel estimation without the need for a ground truth table. This method generated inputs by adding noise to the received signal, while the noisy signal served as the data label. Although this approach yielded notable results, it faced limitations in real-world scenarios where accurate noise variance estimates were often unavailable. Although the techniques mentioned above [15, 11, 12, 13, 14, 16] demonstrate satisfactory performance, they still encounter challenges related to scalability and estimation accuracy, and rely on unrealistic dataset assumptions that may not adequately reflect real-world scenarios.
A significant limitation of the data-driven techniques, as mentioned earlier, was that most proposed schemes were trained and evaluated on small-scale systems, typically involving only tens of RIS and antenna elements. Although these schemes perform effectively in moderate-scale scenarios, their scalability becomes a critical limitation when applied to large-scale systems involving hundreds of RIS elements and thousands of antennas, rendering them impractical for RIS-enabled XL-MIMO deployments [5]. Training such large systems requires complex and high-end infrastructure to handle the computational demands, posing a significant challenge to the feasibility of these approaches. However, several studies [5, 6, 10] have highlighted the presence of spatial correlation in the cascaded channel. In particular, some works have conducted a detailed analysis of spatially correlated RIS with antenna elements, providing valuable insights into the impact of correlation on system performance [18, 19, 20]. Therefore, this paper leverages the spatial correlation inherent in the cascaded channel to reduce model complexity, enabling efficient operation in XL-MIMO systems on resource-constrained devices. To effectively train our model, we employ an autoencoder-based architecture. The work presented in [23] has used an autoencoder-based architecture (UNet) combined with ResNet to capture spatial correlations between inputs and outputs, achieving strong performance in mmWave massive MIMO systems.
I-B Contributions
This is the first work to address the challenges of cascaded channel estimation for RIS-assisted XL-MIMO systems. More specifically, the key contributions of our work are summarized as follows:
  • We propose a lightweight deep learning framework for efficient cascaded channel estimation in RIS-enabled XL-MIMO systems. The proposed framework is designed with reduced computational complexity, rendering it highly suitable for deployment on resource-constrained edge devices, which are integral to the infrastructure of next-generation wireless communication networks.
  • We leverage the spatial correlations in the channel to propose a patch-based training mechanism, where carefully selected small patches encapsulate the essential features of each data point. This approach effectively reduces the input dimensionality to patch-level representations while preserving essential information, thereby facilitating scalable training for large-scale systems.
  • We propose a UNet 3+-based architecture, customized with a dedicated denoising module to effectively extract noise from the channel coefficients and enable efficient cascaded channel estimation. The proposed architecture demonstrates reliable performance even when trained using limited data, maintaining high accuracy with small patch sizes. This highlights its robustness, computational efficiency, and adaptability for deployment on low-resource edge devices.
  • We generate a synthetic dataset based on realistic real-world settings to accurately capture the spatial characteristics of wireless communication environments, moving beyond purely random channel modeling for improved fidelity and reliability.
  • Extensive simulations are conducted under diverse environmental conditions, and the results show that the proposed solution maintains consistent complexity regardless of the size of the RIS-enabled XL-MIMO system, demonstrating its scalability and suitability for edge device deployment.
The remainder of the paper is organized as follows. Section II introduces the system model and the proposed channel estimation framework. Section III provides a detailed description of the proposed model architecture. Section IV discusses the spatial correlation between antennas and RIS elements. Section V presents the patch-based channel estimation approach along with an in-depth effectiveness analysis. Section VI analyzes the computational complexity of the proposed method. Section VII outlines the simulation settings and presents the results. Finally, Section VIII concludes the paper with key findings.
Notation: $T$ and $H$ represent the transpose and Hermitian transpose of a matrix, while $\mathbb{R}$ and $\mathbb{C}$ denote the sets of real and complex numbers, respectively. The function $\text{diag}(\cdot)$ constructs a diagonal matrix, and $\|\cdot\|_{F}$ denotes the Frobenius norm, and $e$ indicates the exponential function. ${\mathbf{I}_{M}}$ represents the identity matrix of size $M\times M$. The operator $(\cdot)^{-1}$ denotes the matrix inverse, while $\mathbb{E}(\cdot)$ indicates the statistical expectation. The symbol $\otimes$ is used to represent the Kronecker product.
II System Model
This paper considers an XL-MIMO system assisted by an RIS panel, for example deployed on the facade of a building, as shown in Fig. 1, to improve the propagation environment. The proposed system model includes a BS equipped with $M$ antennas arranged in a uniform planar array configuration, where $M_{h}$ and $M_{v}$ are the numbers of horizontal and vertical antennas, communicating with $K$ single-antenna users. Moreover, the RIS, with passive reflecting elements $N=N_{h}\times N_{v}$, is arranged in a uniform planar array, where $N_{h}$ and $N_{v}$ are the numbers of horizontal and vertical elements, respectively. Additionally, each RIS element acts as a single point source, aggregating incoming signals and retransmitting them to the BS.
Figure 1: The uplink of the RIS-assisted XL-MIMO system.
We consider a real-world scenario where a signal propagates through a three-dimensional (3D) space, with a local spherical coordinate system defined such that $\phi$ represents the azimuth angle and $\psi$ denotes the elevation angle [18]. The reflecting and antenna elements are arranged in a two-dimensional grid within the YZ plane. Each element has $d_{H}$ and $d_{V}$ corresponding to the horizontal and vertical spacing, respectively. To systematically index the elements, they are ordered row by row and denoted as $n\in[1,N]$ RIS elements and $m\in[1,M]$ antenna elements. The position of the $n$-th element relative to the origin is:
$\boldsymbol{d}_{n}=\begin{bmatrix}0,\alpha_{n}d_{H},\gamma_{n}d_{V}\end{bmatrix}^{T},$ (1)
where the horizontal $\alpha_{n}$ and vertical $\gamma_{n}$ indices of the $n$-th element are represented as:
$\alpha_{n}=\text{mod}(n-1,N_{h}),\quad\gamma_{n}=\lfloor(n-1)/N_{h}\rfloor,$ (2)
[…]
$\mathbf{a}(\phi,\psi)=\begin{bmatrix}e^{j\boldsymbol{\omega}(\phi,\psi)^{T}\boldsymbol{d}_{1}},\dots,e^{j\boldsymbol{\omega}(\phi,\psi)^{T}\boldsymbol{d}_{N}}\end{bmatrix}^{T},$ (3)
where $\boldsymbol{\omega}(\phi,\psi)\in\mathbb{R}^{3\times 1}$ denotes the wave vector, defined as:
$\boldsymbol{\omega}(\phi,\psi)=\frac{2\pi}{\lambda}\begin{bmatrix}\cos(\psi)\cos(\phi)\\ \cos(\psi)\sin(\phi)\\ \sin(\psi)\end{bmatrix},$ (4)
Furthermore, a Line of Sight (LoS) is assumed to exist between the users and the RIS and between the RIS and the BS. We define the channel between the $k$-th user and the RIS elements as a Rician fading channel represented by $\mathbf{f}_{k}\in\mathbb{C}^{N\times 1}$, which can be expressed as follows:
$\mathbf{f}_{k}=\sqrt{\frac{\eta_{r}}{\eta_{r}+1}}\mathbf{f}^{{LoS}}_{k}+\sqrt{\frac{1}{\eta_{r}+1}}\mathbf{f}^{{NLoS}}_{k},$ (5)
Moreover, the LoS channel component $\mathbf{f}^{LoS}_{k}$ between the user and RIS is given by:
$\mathbf{f}^{LoS}_{k}=\mathbf{a}_{r}^{H}(\phi_{r,k},\psi_{r,k}),$ (6)
where $\phi_{r,k}$ and $\psi_{r,k}$ represent the azimuth and elevation angles of the $k$-th user’s signal at the RIS. For a uniform planar array with $N=N_{h}\times N_{v}$ reflecting elements, the array response vector $\mathbf{a}_{r}(\phi,\psi)\in\mathbb{C}^{N\times 1}$ can be modeled as:
$\begin{split}\!\!\mathbf{a}_{r}(\phi,\psi)\!\!=\!\!\frac{1}{\sqrt{N}}[1,\ldots,e^{j\frac{2\pi}{\lambda}\left(\alpha_{n}\cos(\psi)\sin(\phi)d_{H}+\gamma_{n}\sin(\phi)d_{V}\right)},\\ \ldots,e^{j\frac{2\pi}{\lambda}\left(\alpha_{N}\cos(\psi)\sin(\phi)d_{H}+\gamma_{N}\sin(\phi)d_{V}\right)}]^{T},\end{split}$ (7)
where $\alpha_{n}$ and $\gamma_{n}$ are the indices of reflecting elements in the horizontal and vertical directions, respectively. We consider the Non-LoS (NLoS) channels as quasi-static Rayleigh fading, which remains constant for the coherence interval. The NLoS channel between the RIS and the user is modeled using correlated Rayleigh fading [25]. The NLoS channel from the RIS to the user is given by:
$\mathbf{f}^{NLoS}_{k}={(\mathbf{R}_{r})^{\frac{1}{2}}}\tilde{\mathbf{f_{k}}},$ (8)
where $\tilde{\mathbf{f_{k}}}\in\mathbb{C}^{N\times 1}$ follows a complex Gaussian distribution $\mathcal{N_{\textbf{c}}}(0,\mathbf{I}_{N})$, and $\mathbf{R}_{r}$ represents the correlation matrix for the RIS discussed in section III. While the Rician channel between the RIS and the BS is denoted by $\mathbf{H}\in\mathbb{C}^{M\times N}$, it can be expressed as follows:
$\mathbf{H}=\sqrt{\frac{\eta_{b}}{\eta_{b}+1}}\mathbf{H}^{{LoS}}+\sqrt{\frac{1}{\eta_{b}+1}}\mathbf{H}^{{NLoS}},$ (9)
where $\mathbf{H}^{\text{LoS}}$ is the Kronecker product of the array response of the BS and the RIS, which can be expressed as follows:
$\mathbf{H}^{LoS}=\mathbf{a}_{b}(\phi_{b},\psi_{b})\otimes\mathbf{a}_{r}^{H}(\phi_{r},\psi_{r}),$ (10)
where $\mathbf{a}_{b}(\phi_{b},\psi_{b})\in\mathbb{C}^{M\times 1}$ and $\mathbf{a}_{r}(\phi_{r},\psi_{r})\in\mathbb{C}^{N\times 1}$ are the array response vectors associated with the RIS and BS, respectively. Here, $\phi_{r}$ is the azimuth angle of departure, $\psi_{r}$ represents the zenith angle of departure of the RIS, and $\phi_{b}$ symbolizes the azimuth angle of arrival, and $\psi_{b}$ represents the zenith angle of arrival of the BS. The array response vector for a uniform planar array at the BS can be expressed similarly to equation (7) as:
$\begin{split}\mathbf{a}_{b}(\phi,\psi)=\frac{1}{\sqrt{M}}[1,\ldots,e^{j\frac{2\pi}{\lambda}\left(\alpha_{m}\cos(\psi)\sin(\phi)d_{H}+\gamma_{m}\sin(\phi)d_{V}\right)},\\ \ldots,e^{j\frac{2\pi}{\lambda}\left(\alpha_{M}\cos(\psi)\sin(\phi)d_{H}+\gamma_{M}\sin(\phi)d_{V}\right)}]^{T},\end{split}$ (11)
where $\alpha_{m}$ and $\gamma_{m}$ are the indices of antenna elements in the horizontal and vertical directions, respectively. The $\mathbf{H}^{NLoS}$ can be expressed as:
$\mathbf{H}^{{NLoS}}={(\mathbf{R}_{b})^{\frac{1}{2}}}\tilde{\mathbf{H}}{(\mathbf{R}_{r})^{\frac{1}{2}}},$ (12)
where $\tilde{\mathbf{H}}$ $\sim$ $\mathcal{N_{\textbf{c}}}(0,\mathbf{I}_{M}\otimes\mathbf{I}_{N})$ and $\mathbf{R}_{\text{b}}$ and $\mathbf{R}_{\text{r}}$ are the correlation matrices for the BS and RIS, respectively. The correlation matrix for the BS is modeled as a Toeplitz matrix using an exponential correlation function $\mathbf{R_{b}}_{[i,j]}=\rho^{|i-j|}$ between $i$-th and $j$-th antenna, where $\rho$ is the correlation factor, which characterizes the correlation between elements based on their relative indices. This exponential model assumes that the correlation decays with the distance between elements. The total 2D correlation matrix $\mathbf{R}_{b}$ is then obtained by taking the Kronecker product of the horizontal and vertical correlation matrices, i.e., $\mathbf{R}_{b}=\mathbf{R}_{M_{h}}\otimes\mathbf{R}_{M_{v}}$, where $\mathbf{R}_{M_{h}}$ and $\mathbf{R}_{M_{v}}$ are the correlation matrices for the horizontal and vertical directions, respectively. Moreover, the direct channel between the $k$-th user and the BS follows a correlated Rayleigh fading model represented by $\mathbf{b}_{k}\in\mathbb{C}^{M\times 1}$, expressed as:
$\mathbf{b}_{k}=(\mathbf{R}_{b})^{\frac{1}{2}}\tilde{\mathbf{b_{k}}},$ (13)
where $\tilde{\mathbf{b_{k}}}$ follows a complex Gaussian distribution $\mathcal{N_{\textbf{c}}}(0,\mathbf{I}_{M})$, and $\mathbf{R}_{b}$ represents the correlation matrix for the BS.
II-A Channel Estimation Framework
The system’s performance is heavily dependent on the availability of precise channel information. We adopted a Time Division Duplex (TDD) protocol, which exploits channel reciprocity to estimate downlink CSI from uplink CSI. The TDD protocol divides communication into time slots, known as coherence intervals, during which the channel response remains approximately constant [25]. We divide the channel estimation framework into two phases: direct channel estimation and cascaded channel estimation. In the first phase, the RIS is turned off, so no reflections occur. The direct channel between the user and the BS is estimated during this stage, as shown in Fig. 2.
Figure 2: Channel estimation framework has two phases: (1) Direct channel estimation with the RIS off, and (2) Cascaded channel estimation with the RIS on using $L\geq N$ distinct phase shift matrices.
Moreover, in the second phase, we estimate the cascaded channel between the user, the RIS, and the BS. Estimating the cascaded channel requires $L$ distinct phase shift matrices, where $L\geq N$, to facilitate accurate channel estimation. For each phase shift matrix, each user is assigned an orthogonal pilot sequence $\mathbf{x}_{k}$ of length $u$, where $u\geq K$, ensuring the pilot’s orthogonality for accurate estimation. For the $k$-th user during the $l$-th subframe, the pilot signal received at the BS, denoted as $\mathbf{y}_{k}\in\mathbb{C}^{M\times 1}$, is given by the following expression:
$\mathbf{y}_{k}=\mathbf{b}_{k}\mathbf{x}_{k}^{H}+\mathbf{H}\text{diag}(\mathbf{\Theta}_{l})\mathbf{f}_{k}\mathbf{x}_{k}^{H}+\mathbf{n}_{k},$ (14)
where $\boldsymbol{\Theta}_{l}=\begin{bmatrix}\beta_{l,1}e^{j\theta_{l,1}},\beta_{l,2}e^{j\theta_{l,2}},\ldots,\beta_{l,N}e^{j\theta_{l,N}}\end{bmatrix}^{T}$ denotes the reflection coefficients of the RIS elements in the $l$-th sub-frame. Here, $0\leq\beta_{l,n}\leq 1$ and $0\leq\theta_{l,n}<2\pi$ indicate the amplitude and phase shift of the $n$-th RIS element, respectively. The noise term $n_{k}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{m})$ represents AWGN. Furthermore, we can thus separate the $k$-th user signal by multiplying the received signal with $\frac{1}{Pu}x_{k}$:
$\mathbf{y}_{k}=\frac{1}{Pu}(\mathbf{b}_{k}\mathbf{x}_{k}^{H}\mathbf{x}_{k}+\mathbf{H}\text{diag}(\mathbf{\Theta}_{l})\mathbf{f}_{k}\mathbf{x}_{k}^{H}\mathbf{x}_{k}+\mathbf{n}_{k}\mathbf{x}_{k}),$ (15)
Since pilot sequences are orthogonal, $\mathbf{x}_{k}^{H}\mathbf{x}_{k}=Pu$, where $P$ and $u$ are the power and length of the pilot, respectively. Here the noise term $\mathbf{v_{k}}=\frac{1}{Pu}\mathbf{x}_{k}\mathbf{n}_{k}$:
$\mathbf{y}_{k}=\mathbf{b}_{k}+\mathbf{H}\text{\text{diag}}(\mathbf{\Theta}_{l})\mathbf{f}_{k}+\mathbf{v}_{k},$ (16)
It is evident that $\text{\text{diag}}(\mathbf{\Theta}_{l})\mathbf{f}_{k}=\text{\text{diag}}(\mathbf{f}_{k})\mathbf{\Theta}_{l}$. For cascaded channel estimation, the process requires received pilot signals across $L$ frames, collected as $\mathbf{Y}_{k}=[\mathbf{y}_{k,1},\mathbf{y}_{k,2},\ldots,\mathbf{y}_{k,L}].$ Similarly, the phase shift matrices across these frames are represented by $\mathbf{S}=[\mathbf{\Theta}_{1},\mathbf{\Theta}_{2},\ldots,\mathbf{\Theta}_{L}]\in\mathbb{C}^{N\times L}.$
$\mathbf{Y}_{k}=\mathbf{b}_{k}+\mathbf{G}_{k}\mathbf{S}+\mathbf{V}_{k},$ (17)
where $\mathbf{G}_{k}=\ \mathbf{H}\,\text{diag}(\mathbf{f}_{k})\in\mathcal{C}^{M\times N}$ represents the cascaded channel and $\mathbf{V}_{k}=[\mathbf{v}_{k,1},\mathbf{v}_{k,2},\ldots,\mathbf{v}_{k,L}]$. According to [14], designing the optimal phase shift matrix $\mathbf{S}$ using a discrete Fourier transform maximizes the received signal power at the BS, which can be expressed as follows:
$\mathbf{S}=\begin{bmatrix}1&W^{1}&\cdots&W^{1.(L-1)}\\ \vdots&\vdots&\ddots&\vdots\\ 1&W^{N}&\cdots&W^{N.(L-1)}\end{bmatrix}\in\mathbb{C}^{N\times L},$ (18)
where $W=e^{j\frac{2\pi}{L}}$ and $\mathbf{S}\mathbf{S}^{H}=L\mathbf{I}_{N}$. Cascaded channel estimation involves recovering $\mathbf{G_{k}}$ from the received pilots. The simplest approach is the LS method [26], which performs effectively when the channel is assumed to be static. The LS representation of the cascaded channel can be expressed as $\mathbf{\tilde{G}}_{k}^{LS}=\mathbf{Y}_{k}\mathbf{S}^{H}(\mathbf{S}\mathbf{S}^{H})^{-1}$. However, treating the channel as static fails to capture its spatial characteristics, especially when compared to data-driven methods. Moreover, under a Rayleigh fading model and assuming the availability of the statistical channel correlation matrix, the optimal estimator is the MMSE estimator. However, when the channel distribution is complex, as in the case of the cascaded channel in our scenario, the MMSE estimator becomes computationally expensive. To reduce this complexity, the MMSE estimator is often approximated by the LMMSE estimator. According to [27], the LMMSE estimator is given by:
$\mathbf{\tilde{G}}_{k}^{\mathrm{LMMSE}}=\mathbf{Y}_{k}(\mathbf{S}^{H}\mathbf{R}_{G_{k}}\mathbf{S}+M\sigma_{V}^{2}\mathbf{I}_{S})^{-1}\mathbf{S}^{H}\mathbf{R}_{G_{k}},$ (19)
where $\mathbf{R}_{G_{k}}=\mathbb{E}(G_{k}^{H}G_{k})$ denotes the statistical channel correlation matrix. However, since $G_{k}$ corresponds to a cascaded channel, it does not follow the Rayleigh fading model. Consequently, the assumptions underlying the LMMSE estimator are violated, leading to a significant performance gap between the LMMSE estimation and the actual channel. To address this limitation, we propose a deep learning-based method that effectively learns the underlying channel structure and mitigates the performance gap introduced by the model mismatch in conventional methods.
III Deep Learning Assisted Channel Estimation
Deep learning-assisted cascaded channel estimation has gained significant popularity in recent years [15, 11, 12, 13, 14, 16]. One of the key reasons for this growing interest is its extremely low inference time, which makes it well-suited for low-latency and URLLC applications. However, existing works primarily focus on small-scale scenarios with only a limited number of antennas and RIS elements. In this paper, we investigate the use of deep learning to estimate the channel in RIS-enabled XL-MIMO systems, which may involve hundreds of RIS elements and thousands of antennas. We leverage the spatial correlation inherent in the cascaded channel, resulting from the closely placed RIS elements and BS antennas. Building on this property, we propose a patch-based training framework that significantly reduces input dimensionality, enabling scalable training for large-scale systems on resource-constrained edge devices. This section outlines the proposed deep learning framework, while Sections IV and V offer in-depth discussions on spatial correlation and the patch-based training strategy.
III-A Theoretical Background
Recent studies have highlighted that autoencoders can learn efficient, low-dimensional data representations by capturing essential patterns and structures while minimizing the reconstruction loss [22, 21, 23]. They also generalize very well, making them effective in modeling complex, high-dimensional distributions. Autoencoders also require less data to train because they learn compact latent representations that retain the most critical features of the input. This latent space not only enables efficient data compression and generalization, but also helps extracting meaningful features, which is particularly beneficial in our case. Despite their effectiveness, autoencoders may encounter reconstruction limitations, as compression into a bottleneck layer can lead to the loss of critical spatial features, hindering the accurate recovery of fine-grained structures. To address this challenge, the authors in [21] proposed the UNet architecture, which mitigates this issue through skip connections that directly transfer high-resolution features from the encoder to the decoder. This design enables UNet to preserve essential spatial correlations while effectively filtering out noise [23].
In our proposed framework, we leverage the UNet 3+ architecture, an advanced extension of UNet, designed to aggregate feature maps from multiple hierarchical levels. UNet 3+ improves upon the standard UNet by incorporating enhanced multi-scale feature aggregation, making it more effective for denoising tasks [24]. While traditional UNet relies on skip connections between corresponding encoder and decoder layers, UNet 3+ introduces full-scale skip connections, which aggregate feature maps from multiple levels instead of just one, as shown in Fig. 3. This results in richer contextual information, allowing the model to capture both fine-grained local details and broader structural patterns more effectively. This enhanced feature aggregation significantly improves denoising performance by preserving critical spatial structures and more effectively suppressing noise, even in scenarios with limited training data. This makes UNet 3+ a better choice for denoising tasks in large-scale systems than the standard UNet.
Figure 3: The proposed UNet 3+ framework employs $l$ levels, each with a dedicated denoising block in the encoder. Filters double, and the resolution halves at each level. Fused encoder and decoder features serve as input to each decoder layer.
III-B Proposed Model Architecture
We propose a UNet 3+ based architecture designed to effectively capture the underlying signal features in the dataset. The framework integrates a dedicated denoising module that employs a subtraction-based structure to enhance channel estimation by isolating the core signal components from additive noise. The architecture, illustrated in Fig. 3, adopts a multi-level encoder–decoder design to exploit spatial hierarchies to break the noise correlation, enabling more effective denoising and improved estimation accuracy. In the encoder, each level in the architecture consists of multi-layer Convolutional Neural Networks (CNNs) with distinct denoising blocks. The output of each higher-level block is downsampled by a factor of two and fed into the subsequent lower-level block, where the number of filters is doubled to capture increasingly complex features. We employ a MaxPooling layer with a (2,2) kernel in the encoder to downsample between levels. In the decoder, the output from lower-level blocks is upsampled to the current level, whereas the feature maps from the corresponding encoder blocks are downsampled to match that level. Additionally, the output from the same level in the encoder is concatenated with these other outputs, enabling effective multi-scale information fusion to preserve spatial structure. Moreover, we utilized bilinear interpolation for upsampling between levels in the network to smoothly increase the resolution of feature maps while maintaining spatial consistency. During the decoder phase, the filter size is reduced by half, which helps to refine the reconstructed output. Finally, an activation function is applied to the aggregated features before generating the final output, introducing nonlinearity and enhancing the model’s ability to learn complex mappings.
III-C Denoising Block Structure
The CNN architecture is well-suited to extract underlying signal features from noisy data effectively. In [11], a subtraction-based mechanism was proposed to eliminate noise from the data, demonstrating effectiveness in denoising. However, this approach was limited to learning models in which the input and output have identical dimensions, a constraint that does not align with our proposed architecture. To address this limitation, the proposed denoising architecture introduces a $1\times 1$ convolutional layer at the beginning of each level in the encoder to align the input and output dimensions, thereby facilitating the application of the denoising mechanism. The denoising block architecture is composed of a CNN with $i$ layers, where each layer consists of a convolution operation, batch normalization, and a non-linear activation function, as illustrated in Fig. 3. The denoising block takes an input and applies a pointwise ($1\times 1$) convolution to align it with the network’s internal dimensions. The transformed input then passes through multiple convolutional layers. Finally, an element-wise subtraction is performed between the $l$-th level’s block input $\mathbf{I}_{l}$ and output $\tilde{\mathbf{z}}_{\theta}(\mathbf{I}_{l})$ to produce the denoised channel matrix, effectively preserving essential features. At the $l$-th level, it can be expressed as:
$\mathbf{Z}_{l}=\mathbf{I}_{l}-\tilde{\mathbf{z}}_{\theta}(\mathbf{I}_{l})$ (20)
where $\tilde{\mathbf{z}}_{\theta}$ denotes the convolutional neural network and $\theta$ represents its parameters. The primary challenge associated with cascaded channel estimation is the presence of additive white Gaussian noise (AWGN) [6]. The simplest estimator for this task is the LS method, as it does not require statistical information. However, channel estimation using LS is often inadequate, as it is highly sensitive to noise, particularly in low Signal-to-Noise Ratio (SNR) conditions. Although it may still provide a rough estimate, its performance significantly degrades in noisy environments. Therefore, we use the LS estimator as input to our model, as it allows for a coarse estimate of the cascaded channel [11]. This initial estimate enables our model to focus on refining the channel representation, rather than learning it from scratch, thereby improving convergence and overall performance. Moreover, to accommodate the complex-valued nature of the data, each sample is decomposed into its real and imaginary components and provided as separate input channels to facilitate effective feature extraction. Thus, the input sample at the first level is represented as $\mathbf{I}_{0}=\left[\text{Re}\left\{\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\right\},\ \text{Im}\left\{\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\right\}\right]$.
IV Spatial Correlation
Spatial correlation refers to the degree of similarity between signals received or transmitted across different antenna elements, primarily influenced by their physical proximity and exposure to similar propagation paths [19]. When the RIS elements and BS antennas are closely spaced, the angular spread of the incoming and outgoing signals becomes narrow, resulting in similar fading patterns across the array. Consequently, the cascaded channel experiences a certain level of spatial correlation [20]. This section provides a detailed discussion on the spatial correlation present in the cascaded channel and explores how it can be effectively leveraged to optimize the training process and reduce computational complexity.
IV-A Spatial Correlation Between RIS Elements
To model spatial correlation in the RIS elements, consider the channel between the $m$-th antenna at the BS and the $N$ elements of the RIS, represented as a complex-valued vector $\mathbf{h}_{m}=[h_{m,1},h_{m,2},\dots,h_{m,N}]\in\mathbb{C}^{N}$. This channel vector can be further expressed as:
$\mathbf{h}_{m}\sim\mathcal{N_{C}}\left(0_{N},[\mathbf{R}_{b}]_{m,m}\mathbf{R}_{r}\right),$ (21)
Let $\mathbf{R}_{r}\in\mathbb{C}^{N\times N}$ represent the correlation matrix of the RIS, which can be expressed as $\mathbf{R}_{r}=\mathbb{E}\left\{\mathbf{h}_{m}\mathbf{h}_{m}^{H}\right\}=\mathbb{E}\left\{\mathbf{a}_{r}(\phi,\theta)\mathbf{a}_{r}(\phi,\theta)^{H}\right\}$ [18]. To determine the correlation between elements, we select another RIS element $n^{\prime}$ and analyze the relationship between their respective channel responses, which can be expressed as:
$\mathbf{R}_{r_{[n,n^{\prime}]}}$ $=\frac{1}{{NM}}\mathbb{E}\Bigg{[}e^{j\frac{2\pi}{\lambda}\big{(}\alpha_{n}\cos(\psi)\sin(\phi)d_{H}+\gamma_{n}\sin(\phi)d_{V}\big{)}}$
$\quad\cdot\left(e^{j\frac{2\pi}{\lambda}\big{(}\alpha_{n^{\prime}}\cos(\psi)\sin(\phi)d_{H}+\gamma_{n^{\prime}}\sin(\phi)d_{V}\big{)}}\right)^{H}\Bigg{]},$ (22)
Multiplying the exponential terms allows for further simplification of the above equation as follows:
$\!\!\mathbf{R}_{r_{[n,n^{\prime}]}}\!\!\!\!\!=\!\!\mathbb{E}\!\left[e^{j\frac{2\pi}{\lambda}\left[(\alpha_{n}\!-\alpha_{n^{\prime}})\!\cos(\psi)\sin(\phi)d_{H}+(\gamma_{n}\!-\gamma_{n^{\prime}})\sin(\phi)d_{V}\right]}\right]\!,\!\!\!\!$ (23)
We consider the case where $\alpha_{n^{\prime}}=\alpha_{n}$, which means that the elements share the same row. This simplifies the expression for $\mathbf{R}_{r_{[n,n^{\prime}]}}$, as the term that involves $N_{h}$ disappears. The expectation can be further expressed using Equation (3) as follows:
[…]
We now proceed to evaluate the integral of $\mathbf{R}_{N}$ with respect to the angular variables $\phi$ and $\psi$:
[…]
When performing the integration over $\phi$, the resulting expression is a classic oscillatory integral that evaluates to the Bessel function of the first kind, $J_{0}$:
$\mathbf{R}_{r_{[n,n^{\prime}]}}=J_{0}\left({\frac{2\pi}{\lambda}||\boldsymbol{d}_{n}-\boldsymbol{d}_{n^{\prime}}||}\right),$ (26)
The correlation function $\mathbf{R}_{r_{[n,n^{\prime}]}}$ follows a zero-order Bessel function of the first kind [20]. This expression captures the correlation between two RIS elements separated by $||\boldsymbol{d}_{n}-\boldsymbol{d}_{n^{\prime}}||$. The Bessel function structure indicates that the correlation oscillates and gradually decays with increasing element separation, reflecting the spatial coherence properties of the scattered waves.
IV-B Spatial Correlation Between RIS and BS
In the previous subsection, we discussed the correlation between the $m$-th antenna and the RIS elements. However, the cascaded channel between the RIS and the BS, denoted as $\mathbf{G}_{k}\in\mathbb{C}^{M\times N}$, consists of $M$ antennas at the BS and $N$ reflecting elements at the RIS. To determine the correlation between the $m^{\prime}$-th antenna and the RIS elements, the corresponding channel coefficients can be expressed as:
$\mathbf{h}_{m^{\prime}}\sim\mathcal{N_{C}}\left(0_{N},[\mathbf{R}_{b}]_{m^{\prime},m^{\prime}}\mathbf{R}_{r}\right),$ (27)
Note that equation (27) follows a structure similar to equation (21), with the key distinction being in the variance terms $[\mathbf{R}_{b}]_{m,m}$ and $[\mathbf{R}_{b}]_{m^{\prime},m^{\prime}}$. These variance terms capture the spatial correlation among antennas, which are not independent but follow the correlation function $[\mathbf{R}_{b}]_{m,m^{\prime}}=\rho^{|m-m^{\prime}|}$, where $\rho$ represents the correlation coefficient. In conclusion, the channel matrix $\mathbf{G}_{k}\in\mathbb{C}^{M\times N}$ exhibits spatial correlation, governed by the correlation structures of RIS and BS. Thus, in this paper, we leverage spatial correlation to minimize training complexity by proposing a patch-based training framework, which is discussed in detail in the following section.
V Patches based Training of Channel State Information
Training machine learning models on edge devices is becoming increasingly popular, driven by the need for low-latency performance and enhanced data privacy [17]. Furthermore, it facilitates the development of robust and adaptive models that are tailored to local conditions, improving the overall performance of edge applications. However, training RIS-enabled XL-MIMO models on edge devices is challenging due to the size of the cascaded channel estimation matrices, $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$, where both $M$ and $N$ can range from hundreds to thousands [4]. This results in massive data sizes, making storing and processing such large amounts of data on resource-constrained edge devices infeasible. To address this challenge, we propose a patch-based localized training mechanism. This method leverages spatial correlation between channels to reduce computational and memory demands, enabling feasible model training on edge devices despite their limited resources.
V-A Patches Based Training Framework
The patch-based training framework addresses the challenges of training large models on resource-constrained devices. In this framework, instead of processing the entire high-resolution data point $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$ at once, the proposed approach divides it into smaller, more manageable patches. Each patch is trained separately, which reduces the memory and computational requirements, making the training process more efficient. Patch-based training is beneficial for applications in edge devices, where computational and storage resources are limited. In the proposed framework, we exploit the spatial correlation present in the cascaded channel to observe that extracting a small number of representative patches from each training example is sufficient for effective model training. These extracted patches from each data point capture the essential features required for training, eliminating the need to process the entire high-dimensional matrix $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$. This is particularly relevant in the context of RIS-assisted XL-MIMO systems, where the number of RIS elements can range from hundreds to thousands. Our proposed approach significantly reduces computational complexity by focusing on smaller patches. This reduction in complexity enables resource-constrained edge devices to perform training at the edge, which would otherwise be infeasible when dealing with large-scale datasets.
As illustrated in Algorithm 1, patches are extracted from each data point in the training set $X=\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}$ along with their corresponding labels $Y={\mathbf{G}}^{(i)}_{k}$ to train the model. The inputs to the algorithm include the training set $X$, the labels $Y$, the patch dimensions $p_{x}$ and $p_{y}$, and the total number of patches total_patches to be generated. Initially, the algorithm calculates the number of patches per data point $p_{\text{dp}}$. The algorithm then iterates over the entire training set and generates $p_{\text{dp}}$ patches for each data point. For each patch, random starting coordinates $x$ and $y$ are selected within the valid range defined by the data dimensions $M$ and $N$. Subsequently, the training samples $X_{dp}$ and labels $Y_{dp}$ are sliced at the selected coordinates to extract patches of size $p_{x}\times p_{y}$ from the data point along with their corresponding labels. These patches are then appended to the lists patches_data and patches_labels. Finally, the algorithm returns patches and labels for use in model training.
Algorithm 1 Patch-Based Training Data Generation
[…]
3:  $p_{\text{dp}}=\frac{\text{total\_patches}}{\text{length}(X)}$
[…]
5:  for  $i=1$ to ${\text{length}(X)}$ do
6:     for  $j=1$ to $p_{\text{dp}}$ do
7:        $x,y\leftarrow\text{rand}(1,N-p_{x}),\text{rand}(1,M-p_{y})$
[…]
10:     end for
11:  end for
[…]
V-B Efficacy of the Proposed Patch-Based Model Training
Using the entire data point $\tilde{\mathbf{G}}^{\text{LS}(i)}_{k}\in\mathbb{C}^{M\times N}$ for model training may not be the most efficient approach, given the inherent spatial correlation between antennas and RIS elements in the received signal at the BS, as discussed in the spatial correlation section [25]. From equations (21) and (27), it is evident that the signals received at the $m$-th and $m^{\prime}$-th antennas of the BS exhibit a significant degree of correlation. Consequently, $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$ share spatial correlation, and training the deep learning model with $\mathbf{h}_{m}$ alone is sufficient. The model can inherently learn the correlation structure of $\mathbf{h}_{m^{\prime}}$, which means that there is no need to explicitly include $\mathbf{h}_{m^{\prime}}$ in the training process.
Deep learning models, especially UNet 3+, are particularly suited for this task because they can capture spatial patterns [22]. In our proposed framework, the encoder compresses the input $\mathbf{h}_{m}$ into a lower-dimensional latent space at the bottleneck, capturing the most important features and correlations in the data. This latent representation inherently encodes the spatial correlation between $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$. The decoder then reconstructs the output from this compact representation. However, since the spatial correlation is implicitly captured in the latent space, the decoder does not need direct access to $\mathbf{h}_{m^{\prime}}$ to understand its structure. The model can generalize and learn the dependencies, making it unnecessary to include both $\mathbf{h}_{m}$ and $\mathbf{h}_{m^{\prime}}$ during training. This reduces the complexity of the model while still allowing it to learn the underlying correlation pattern, ensuring that the model performs effectively even with extremely large-scale RIS-based MIMO systems.
VI Computational Complexity
The computational complexity of the proposed method mainly depends on the encoder’s complexity and the decoder’s architecture. The complexity of the encoder in the proposed UNet 3+ based architecture is primarily governed by the convolutional operations and the downsampling (pooling) steps. At each encoder level, two convolutional layers are applied, followed by a max-pooling operation. Given an input feature map of dimensions $H\times W\times C^{\text{in}}$, a convolutional layer with $C^{\text{out}}$ filters and a kernel size of $K\times K$ incurs a computational complexity of $\mathcal{O}(HWC^{\text{in}}C^{\text{out}}K^{2})$. Since each level in the encoder of the proposed architecture comprises two convolutional layers, along with a $1\times 1$ convolution for dimensional alignment and a max-pooling operation, the total computational complexity at level $l$, denoted by $\mathcal{O}_{l}^{e}$, is given by:
$\mathcal{O}_{l}^{e}=\mathcal{O}(2H_{l}W_{l}C^{\text{in}}_{l}C^{\text{out}}_{l}K^{2}+2H_{l}W_{l}C^{\text{out}}_{l})$ (28)
Given that the term $2H_{l}W_{l}C^{\text{out}}_{l}$ contributes relatively little to the overall complexity, it can be considered negligible in the computational cost analysis. Furthermore, the max-pooling operation reduces the spatial dimensions by a factor of two at each level, while the number of feature channels doubles. Thus, for level $l$, the dimensions of the feature map evolve as follows:
$H_{l}=\frac{H_{0}}{2^{l}},\quad W_{l}=\frac{W_{0}}{2^{l}},\quad C^{in}_{l}=2^{l}C_{0}^{out},$ (29)
where $H_{0}\times W_{0}$ represents the spatial dimensions of the input feature map. By summing over all levels from $l=0$ to $L-1$, the total computational complexity of the encoder, represented as $\mathcal{O}^{e}$, can be expressed as:
$\mathcal{O}^{e}=\mathcal{O}\left(\sum_{l=0}^{L-1}2\frac{H_{0}W_{0}}{4^{l}}C^{in}_{l}(2C^{in}_{l})K^{2}\right),$ (30)
Substituting $C^{in}_{l}=2^{l}C_{0}^{out}$, the complexity simplifies to:
$\mathcal{O}^{e}$ $=\mathcal{O}\left(4H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}\sum_{l=0}^{L-1}\frac{4^{l}}{4^{l}}\right)$
$\approx\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L\right)$ (31)
Eq. (31) highlights that the computational cost of the encoder scales linearly with the spatial resolution $H_{0}\times W_{0}$, and quadratically with the number of output filters $\left(C_{0}^{\text{out}}\right)$. Furthermore, deeper networks result in increased computational overhead due to the growing number of feature maps. Moreover, the complexity of the bottleneck layer adopts the same convolutional structure, denoted by $\mathcal{O}^{b}$, is given by:
$\mathcal{O}^{b}=\mathcal{O}\left(H_{L}W_{L}(C^{in}_{l})^{2}K^{2}\right)\approx{H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}},$ (32)
The computational complexity of the decoder in the UNet3+ architecture follows a pattern similar to the encoder, but in reverse. At each decoder level, feature maps are upsampled/downsampled and concatenated before being processed by convolutional layers. The number of output channels decreases as we move towards the output layer, the complexity of $l$-th decoder layer $\mathcal{O}_{l}^{d}$ expressed as :
$\mathcal{O}_{l}^{d}=\mathcal{O}(H_{l}W_{l}C_{\text{concat}}C^{in}_{l}K^{2}),$ (33)
In the decoder, each level processes feature maps of size $H_{l}\times W_{l}$ with $C^{in}_{l}$ channels. Furthermore, the concatenation operation aggregates features from multiple levels, increasing the channel dimension. As a result, the total number of channels after concatenation becomes $C_{\text{concat}}=L\times C^{in}_{l}$. Consequently, the total decoder complexity $\mathcal{O}^{d}$ is given by:
$\mathcal{O}^{d}=\mathcal{O}\left(\sum_{l=0}^{L-1}\frac{H_{0}W_{0}}{4^{l}}L(C_{l}^{in})^{2}K^{2}\right),$ (34)
Substituting $C^{in}_{l}=2^{l}C_{0}^{out}$, the complexity simplifies to:
$\mathcal{O}^{d}=\mathcal{O}\left(H_{0}W_{0}(C_{0}^{out})^{2}K^{2}L^{2}\right),$ (35)
Overall, the computational complexity can be expressed as the sum of the complexities of the encoder, bottleneck, and decoder. Thus, the total complexity $\mathcal{O}_{\text{total}}$ becomes:
$\mathcal{O}_{\text{total}}=$ $\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L\right)+{H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}}$
$+\mathcal{O}\left(H_{0}W_{0}(C_{0}^{\text{out}})^{2}K^{2}L^{2}\right),$ (36)
From (VI), it is evident that the overall computational complexity is directly proportional to the size of the input feature map $H_{0}\times W_{0}$. This implies that reducing the spatial dimensions of the input features can significantly reduce the computational cost of the proposed model.
VII Simulation
This section evaluates the performance of the proposed method. The simulation setup consists of 1024 antennas arranged as $M_{h}=32$ and $M_{v}=32$ planar array, and 128 RIS elements arranged as $N_{h}=16$ and $N_{v}=8$ planar array. The model uses 32 base filters, which double with each depth level in the encoder and are halved progressively in the decoder. Moreover, ReLU activation functions are applied throughout, with a linear function at the output layer. Training begins with an initial learning rate of 0.004, managed by an exponential decay scheduler and optimized via the Adam algorithm. Network configurations include a depth of 3 levels, with 2 convolutional layers per level, a batch size of 32, and 40 training epochs.
After reviewing the related literature, we identified [11] and [14] as the most relevant to our work. While these studies have made a satisfactory contribution to cascaded channel estimation, they are limited by using a small number of antennas and elements. Specifically, their approach involves processing large matrices $\tilde{\mathbf{G}}^{(i)}_{k}\in\mathbb{C}^{M\times N}$ for each data point and training on a dataset containing over 100,000 samples. Given the scale and complexity of XL-MIMO systems, where the number of antennas and RIS elements can be in the order of hundreds or even thousands, extending these approaches to such large-scale systems is not a feasible option due to their computational and memory demands. Consequently, we determined that a fair comparison between our proposed method and these works would not be feasible.
Moreover, we conducted an extensive literature review and found no existing work on training RIS-based XL-MIMO systems. However, we found relevant works mentioned in [26], [27], and [7], which can be used for comparison with our results. Therefore, in this paper, we compare our proposed method with the LS method [26], LMMSE method [27], and B-LMMSE method [7] to evaluate the significance of our results. In the B-LMMSE method, we utilize the binary reflection control strategy proposed in [7], and subsequently apply the LMMSE estimator for channel estimation. The performance of the proposed scheme is measured in terms of the normalized mean squared error (NMSE) $=\frac{E(\|G-\hat{G}_{\text{proposed}}\|_{F}^{2})}{E(\|G\|_{F}^{2})},$ where $\hat{G}_{\text{proposed}}$ and $G$ represent the channel estimated by the proposed method and the ground truth channel, respectively.
In this paper, we generated a synthetic dataset based on the system models described in Section III. The dataset was created while considering the memory constraints of edge devices. Specifically, a total of 10,000 data samples were generated, and from each sample, a representative data patch was extracted to encapsulate the underlying characteristics of the corresponding data point. As the antenna array has dimensions of $32\times 32$ and the RIS elements are arranged as $16\times 8$, the patch sizes were selected as $16\times 32$ and $32\times 32$ to align with the spatial correlation properties between the antenna and RIS elements. The dataset was split into training and validation sets, with 70% allocated for training and 30% for validation. Additionally, we created a separate testing dataset for different SNRs, with a sample size of 2,000 for each SNR value. During the training phase, the model was provided with a representative patch extracted from each data sample to reduce the input size, thereby addressing the computational constraints of resource-limited edge devices. In contrast, full-resolution samples ($\tilde{\mathbf{G}}^{(i)}_{k}\in\mathbb{C}^{M\times N}$) were used during testing to evaluate the model’s performance.
TABLE I: Simulation parameters
Parameter Value Parameter Value
Antennas (M) 1024 RIS Elements (N) 128
$M_{h}$ 32 $M_{v}$ 32
$N_{h}$ 16 $N_{v}$ 8
$\eta_{r}$ 10 $\eta_{b}$ 10
$\rho$ 0.8 Carrier Frequency 7.8 Ghz
Base Filters (BF) 32 $n_{th}$ Level Filters $BF\times 2^{l}$
Activation Function Relu Last Activation Linear
Learning Rate (LR) 0.004 LR Scheduler Exp. Decay
Levels 3 Optimizer Adam
Batch Size 32 Epochs 40
Figure 4: Performance comparison between a $16\times 32$ patch size and the results obtained with a $32\times 32$ patch size after doubling the RIS dimensions to $32$.
Fig. 4 compares the NMSE performance of the proposed method using different patch shapes with that of the B-LMMSE, LMMSE, and LS methods. Each point in the figure represents the average NMSE over 2,000 test samples. The results demonstrate that the proposed architecture significantly outperforms the LS method, particularly at low SNR values, where it exhibits substantially better accuracy. Moreover, B-LMMSE and LMMSE initially perform similarly to the proposed method, but the performance gap becomes more significant as the SNR increases toward 0 dB. Beyond this point, the LMMSE method begins to follow the performance trend of LS, while the B-LMMSE performs worse than LS due to the binary reflection strategy, where only one element is active at a time, resulting in insufficient received power for accurate channel estimation. Additionally, a slight variation in NMSE is observed between different patch sizes. Specifically, the $16\times 32$ patch consistently achieves higher NMSE compared to the $32\times 32$ patch. This improvement stems from the increased RIS dimensions, which enable slightly better extraction of spatial correlations between the RIS elements and BS antennas. The NMSE variation between the above patch shapes ranges from approximately 2% to 15%. Notably, the difference is more pronounced at low SNR levels, where the impact of noise is higher, and it gradually decreases as the SNR increases.
Figure 5: Compared performance by breaking each data point into patches of size $32\times 32$ for testing and evaluating it against using the entire data point without patching.
Fig. 5 presents a comparative analysis of training performance with and without patches during the testing phase. The motivation for employing patches in testing is to evaluate the prediction performance on patches of the same size as those used in training $32\times 32$ versus prediction on the data point of size $1024\times 128$. This allows us to assess the impact of patch-based training when applied to patch-based versus full-resolution testing. As observed in Fig. 5, there is a slight improvement in NMSE when testing with patches compared to testing without patches. The improvement ranges from 0.1% to 8%, with the highest improvement occurring at low SNR values. Specifically, a peak improvement of 8% is observed at $-5$ dB SNR. This result can be attributed to training on patches, enabling the model to estimate noise more effectively at a finer level in low SNR conditions. However, while testing with patches provides a slight performance improvement over testing without patches, it comes at the cost of significantly increased computational time. Specifically, testing with patches requires approximately four times more computation, which may pose challenges in low-latency scenarios. This additional time overhead stems from dividing data into patches, performing individual predictions, and recombining them to estimate the channel.
Figure 6: Performance comparison of the direct channel estimation from the user to BS.
Fig. 6 presents a performance comparison of the direct channel estimation from the user to BS. Similar to the cascaded channel training, we consider 10,000 samples of the direct link channel, denoted as $\mathbf{b}_{k}\in\mathcal{C}^{M\times 1}$. To make the data suitable for training, we transform this 1D data into a 2D representation of size $32\times 32$, leveraging that these dimensions correspond to the antennas’ horizontal and vertical arrangement. As observed in Fig. 6, there is a significant difference in NMSE between the proposed method and the LS algorithm, particularly at low SNR values. This discrepancy arises due to the high noise levels in low SNR scenarios, which pose a major challenge for LS. However, the proposed algorithm demonstrates superior performance, providing reliable channel estimates even under severe noise conditions. Furthermore, the performance gap between the proposed method and LS decreases as SNR increases. This trend can be attributed to the lower noise levels at higher SNR values, which reduce the estimation error for both methods.
Figure 7: Comparative analysis between correlated and uncorrelated fading assumptions for NLoS channels, along with a comparison of results using different patch sizes for each case.
Fig. 7 compares the NMSE performance of correlated and uncorrelated Rayleigh fading channels for different training patch sizes. The correlated channel exhibits the lowest NMSE compared to the uncorrelated Rayleigh fading channel. This is because correlated fading patterns are more structured, allowing deep learning models to capture better and exploit these patterns. However, in practical scenarios, achieving uncorrelated Rayleigh fading is nearly impossible, as there are always dominant directions that carry most of the signal energy. Furthermore, at low SNR levels, the NMSE performance of both correlated and uncorrelated channels is similar due to the overwhelming presence of noise. A noticeable performance difference between correlated and uncorrelated Rayleigh fading channels begins to emerge at an SNR of 0 dB, with an initial gap of approximately 0.5 dB. As the SNR increases, this gap widens significantly, with the correlated channel demonstrating up to a 2.5 dB improvement at an SNR of 15 dB.
Figure 8: Performance comparison between LS and the proposed mechanism by varying the pilot overhead through changes in the phase shift matrix $L$ at an SNR of 5 dB.
Fig. 8 illustrates the impact of increasing the number of phase shift matrices (pilot overhead) on NMSE performance, comparing it with LS and LMMSE methods. As observed, increasing the phase shift matrix positively influences NMSE by reducing estimation error. However, introducing excessive phase shifts leads to significant pilot overhead, which is undesirable in practical communication systems. A sharp decrease in NMSE is noticeable when transitioning from $L=96$ to $L=128$. This improvement is because, as $L$ approaches 128, the system moves closer to achieving a full-rank LS matrix. A full-rank LS matrix ensures better-conditioned channel estimation, leading to effective model training. Beyond $L=128$, NMSE continues to decrease, but at a slower rate, as the benefits of increasing $L$ diminish. Moreover, our proposed method yields good results, while the difference between the LS and LMMSE methods remains minimal, as both rely on the full-rank assumption.
Figure 9: Comparative evaluation of varying the number of RIS $N$ and antennas $M$ on system performance at an SNR of 10 dB.
Fig. 9 illustrates the impact of varying the number of BS antennas $M$ and RIS elements $N$ on NMSE performance compared to LS estimation. As shown, increasing $M$ has a negligible effect on NMSE, even when training with different patch sizes. This is because the training process relies on patches, allowing the model to learn latent representation in the data irrespective of the input size. Consequently, during inference, varying $M$ does not significantly impact the results. A similar trend is observed for N, where the number of RIS elements $N$ is increased. This indicates that training the model with patches enables it to capture underlying data correlations independent of $M$ and $N$. In contrast, traditional schemes are inherently affected by variations in these parameters, highlighting a key advantage of the proposed learning-based approach.
VIII Conclusions
In this work, we addressed the critical challenge of cascaded channel estimation in XL-MIMO systems empowered by RIS. Recognizing the limitations of traditional and existing deep learning approaches in resource-constrained environments, we proposed a novel, lightweight framework tailored for next-generation wireless networks. Using spatial correlation and patch-based training, our solution effectively reduces the dimensionality of the training data while preserving key features. We further enhance performance using a customized UNet 3+ architecture with integrated denoising capabilities, allowing accurate estimation even with limited training data. Extensive evaluations confirm that our framework enhances estimation accuracy and significantly reduces computational overhead, maintaining efficiency even as the number of antennas and RIS elements scales in XL-MIMO systems, making it well-suited for real-time deployment on edge devices. This pioneering approach lays the foundation for scalable, energy-efficient, and practical RIS-enabled XL-MIMO systems in 6G and beyond.